# Pluricanonical Geometry of Varieties Isogenous to a Product: Chevalley-Weil Theory and Pluricanonical Decompositions of Abelian Covers

### v1

@@ 3.1. The Chevalley-Weil Formula for Pluricanonical Representations
Let $C$ be a compact Riemann surface and $G$ a finite group with a faithful action
@@
Theorem 3.4 (Chevalley-Weil Formula for Pluricanonical Representations).
@@
Moreover, every $p\in Fix(C)$ maps to a branch point of $F$. Since $h_{i}$ is the local monodromy, we have
@@
${\displaystyle\frac{1}{|G|}\sum_{p\in Fix(C)}\sum_{\begin{subarray}{c}g\in G_{p}\\ g\neq 1_{G}\end{subarray}}\chi(g)\frac{J_{p}(g)^{\nu_{g}}}{1-J_{p}(g)}}=\frac{1}{|G|}\sum_{i=1}^{r}\frac{|G|}{n_{i}}\sum_{l=1}^{n_{i}-1}\chi(h_{i}^{l})\frac{\zeta_{n_{i}}^{l\nu_{h_{i}^{l}}}}{1-\zeta_{n_{i}}^{l}}=\sum_{i=1}^{r}\frac{1}{n_{i}}\sum_{l=1}^{n_{i}-1}\frac{\sum_{\alpha=0}^{n_{i}-1}N_{i,\alpha}\cdot\zeta_{n_{i}}^{l(\alpha+\nu_{h_{i}^{l}})}}{1-\zeta_{n_{i}}^{l}}.$
@@
We want to point out that in oder to determine the integers $N_{i,\alpha}$ it is not necessary to know the representation $\varrho\colon G\to\operatorname{GL}(V)$. Indeed, it suffices to know the character $\chi\in\operatorname{Irr}(G)$ of $\varrho$, as explained in the "Computational Remark 2.9" of [FG16].
@@ 4. Pluricanonical Systems of an Abelian Cover
Let $X\cong(C_{1}\times\cdots\times C_{n})/G$ be a variety isogenous to a product of unmixed type. Consider the associated cover (see diagram (1))
@@
In the preceding section, we described the computation of the character $\chi_{m}$ of the representation
@@
Let $\pi\colon X\to Y$ be a finite abelian cover between complex manifolds with Galois group $G$. Let $m$ be a non negative integer and $\chi\in\operatorname{Irr}(G)$ a character. Our goal is to compute the eigensheaves
$(\pi_{\ast}K_{X}^{\otimes m})^{\chi}(U):=\{\omega\in K_{X}^{\otimes m}(\pi^{-1}(U))\ \mid\ g^{\ast}\omega=\chi(g)\cdot\omega\quad\makebox{for all}\quad g\in G\}.$
@@
$a_{j}(x)t^{k-m}\frac{(dx_{1}\wedge\ldots\wedge dx_{n})^{\otimes m}}{t^{(d-1)m}}=\psi_{ij}^{\ast}\bigg(a_{i}(x)t^{k-m}\frac{(dx_{1}\wedge\ldots\wedge dx_{n})^{\otimes m}}{t^{(d-1)m}}\bigg)=a_{i}(x)g_{ij}^{k-dm}t^{k-m}\frac{(dx_{1}\wedge\ldots\wedge dx_{n})^{\otimes m}}{t^{(d-1)m}}.$
This implies $a_{i}(x)=g_{ij}^{dm-k}a_{j}(x)$. In other words, the collection $\{a_{i}(x)\}_{i}$ defines a section of the line bundle $L^{\otimes(dm-k)}=\mathcal{O}_{Y}(mB)\otimes L^{\otimes(-k)}$. It follows that the map
@@
(4) $L_{\chi_{k}}\cong L^{\otimes k},\qquad 0\leq k\leq d-1.$
@@
$\pi_{\ast}(K_{X}^{{\otimes m}})^{\chi_{k}}=\pi_{\ast}(K_{X}^{{\otimes m}})^{\chi_{(k+ld)}}=L^{\otimes(dm-k-ld)}\otimes K_{Y}^{{\otimes m}}=L^{\otimes(dm-k-ld-d+k)}\otimes L^{\otimes(d-k)}\otimes K_{Y}^{{\otimes m}}.$
If $k\neq 0$, by (4) of Remark 4.4 the previous chain of equalities amounts to
@@
Then the formula follows from Remark 4.4.
Let us now deal with the general case. Given a finite abelian cover $\pi\colon X\to Y$ between complex manifolds with group $G$, we denote by $R$ the ramification divisor and by $D$ the branch divisor. If $T$ is an irreducible component of $R$ and $p\in T$ a general point, we denote by $H$ the stabilizer group of $p$, which in fact does not depend on the choice of $p$, and hence it is also called the stabilizer of $T$. Then, by Cartan’s Lemma [Cartan, Lemma 1], the tangent representation at $p$ is faithful and looks as follows
@@
(5) $p^{\ast}(\mathcal{O}_{Y}(D_{(H,\psi)}))=\mathcal{O}_{Z}(B)\qquad\makebox{and}\qquad p^{\ast}K_{Y}^{\otimes m}=K_{Z}^{\otimes m}.$
@@
(6) $p_{\ast}M^{\otimes(-k)}=\bigoplus_{\underset{\eta_{|H}=\psi^{k}}{\eta\in\operatorname{Irr}(G)}}L^{-1}_{\eta}.$
@@
(7) $(\pi_{\ast}K_{X}^{\otimes m})^{\chi}=(p_{\ast}g_{\ast}K_{X}^{\otimes m})^{\chi}\cong\left(\bigoplus_{k=0}^{d-1}p_{\ast}\left(\mathcal{O}_{Z}\left(\mu_{k}B\right)\otimes M_{\psi^{k}}^{-1}\otimes K_{Z}^{\otimes m}\right)\right)^{\chi}$
@@
There exists a unique $0\leq k_{0}\leq d-1$ such that $\chi_{|H}=\psi^{k_{0}}$. Hence, only the $k_{0}$-summand survives in (7) and by (5) the line bundle $(\pi_{\ast}K_{X}^{\otimes m})^{\chi}$ is isomorphic to
$\left(p_{\ast}\left(p^{\ast}\left(\mathcal{O}_{Y}\left(\mu_{k_{0}}D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\right)\otimes M_{\psi^{k_{0}}}^{-1}\right)\right)^{\chi}\cong\mathcal{O}_{Y}\left(\mu_{k_{0}}D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\otimes\left(p_{\ast}M_{\psi^{k_{0}}}^{-1}\right)^{\chi},$
where the last isomorphism is given by projection formula. Finally, by (6) the previous line bundle is isomorphic to
$\mathcal{O}_{Y}\left(\mu_{k_{0}}D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\otimes\left(\bigoplus_{\underset{\eta_{|H}=\psi^{k_{0}}}{\eta\in\operatorname{Irr}(G)}}L^{-1}_{\eta}\right)^{\chi}\cong\mathcal{O}_{Y}\left(\mu_{k_{0}}D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\otimes L^{-1}_{\chi}.$
Hence, following the maps, we get the isomorphism of line bundles
@@
$r:=k_{0}-m+\left\lceil\frac{m-k_{0}}{d}\right\rceil d.$
@@
(8) $\pi_{\ast}(K_{X}^{\otimes m})^{\chi}\cong\mathcal{O}_{Y}\left(\sum_{(H,\psi)}\mu(m,H,\psi,\chi)\ D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\otimes L_{\chi}^{-1},$
@@
(9) $H^{0}(X,K_{X}^{\otimes m})=\bigoplus_{\chi\in\operatorname{Irr}(G)}\prod_{(H,\psi)}t_{(H,\psi)}^{r(m,H,\psi,\chi)}\ \pi^{\ast}H^{0}\left(Y,\mathcal{O}_{Y}\left(\sum_{(H,\psi)}\mu(m,H,\psi,\chi)\ D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\otimes L_{\chi}^{-1}\right),$
@@
Since $\pi_{\ast}(K_{X}^{\otimes m})$ is reflexiv by [Har80, Corollary 1.7], it is enough to prove the result when $X$ is smooth. Our discussion above shows that the line bundle isomorphism in (8) holds in codimension 1. Hence, it holds globally by Hartog’s extension theorem and the proof follows immediately. ∎
@@
Formula (8) in Theorem 4.7 has been derived by Pardini [Par91, Proposition 4.1, c)] for $m=1$ and by Bauer-Pignatelli [BP21, Proposition 2.5] for $m=2$. They write them as follows
@@
In [Lie03] Liedtke derived formula (9) in Theorem 4.7 for $m=1$ and wrote it as follows:
@@
(10) $L_{\chi^{-1}}=L_{\chi}^{-1}\otimes\mathcal{O}_{Y}\bigg(\sum_{\underset{\chi_{|H}\neq\psi^{0}}{(H,\psi)}}D_{(H,\psi)}\bigg).$
Hence, by using (10) together with (8), we obtain
@@
whence we obtain Bauer-Pignatelli’s result. Assume now $m=1$. By using (10) together with (9), where we replace $\chi^{-1}$ with $\chi$, we get
@@ 4.1. Pluricanonical Systems of an Unmixed VIP with Abelian Group
Let $X\cong(C_{1}\times\dots\times C_{n})/G$ be a variety isogenous to a product of unmixed type with abelian group $G$. Recall that we have the diagram (1):
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${C_{1}\times\dots\times C_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f_{1},\dots,f_{n})}$$\scriptstyle{\Xi}$${C_{1}^{\prime}\times\dots\times C_{n}^{\prime}}$
@@
In summary, this provides a method to compute each summand in the decomposition (11).
@@
Consequently, the integers $r(m,H,\psi,\chi)$ and $\mu(m,H,\psi,\chi)$ are determined immediately. The only remaining data needed to apply Theorem 4.7 are the line bundles $L_{\chi}$, for which the following formula holds [Par91, (2.14) of Proposition 2.1]:
@@
where $d_{\chi}:=\operatorname{ord}(\chi)$. Thus, we have a concrete procedure to compute $L_{\chi}^{\otimes d_{\chi}}$, which determines the line bundle $L_{\chi}$ up to torsion. Only in the special case where all curves $C_{i}^{\prime}=\mathbb{P}^{1}$ we can immediately recover $L_{\chi}$ itself.
5. Pluricanonical maps of a VIP
In this section, we analyze the $m$-canonical map of a VIP. Our main goal is to study its birationality for a given $m\geq 1$. Since such a variety is of general type, there exists a positive integer $m_{0}$ such that this map is birational for all $m\geq m_{0}$.
In the first subsection, we collect some general observations on the $m$-canonical maps of a VIP.
In the second subsection, we turn to representation-theoretic methods. More precisely, we consider an arbitrary abelian cover $\pi\colon X\to Y$ and present two criteria for the birationality of

### v2

@@ 2.1. The Chevalley-Weil Formula for Pluricanonical Representations
Let $C$ be a compact Riemann surface and $G$ be a finite group with a faithful action
@@
Theorem 2.4 (Chevalley-Weil Formula for Pluricanonical Representations, cf. [CheWeil34, Weil35]).
@@
Moreover, every $p\in\operatorname{Fix}(C)$ maps to a branch point of $F$. Since $h_{i}$ is the local monodromy, we have
@@
${\displaystyle\frac{1}{|G|}\sum_{p\in\operatorname{Fix}(C)}\sum_{\begin{subarray}{c}g\in G_{p}\\ g\neq 1_{G}\end{subarray}}\chi(g)\frac{J_{p}(g)^{\nu_{g}}}{1-J_{p}(g)}}=\frac{1}{|G|}\sum_{i=1}^{r}\frac{|G|}{n_{i}}\sum_{l=1}^{n_{i}-1}\chi(h_{i}^{l})\frac{\zeta_{n_{i}}^{l\nu_{h_{i}^{l}}}}{1-\zeta_{n_{i}}^{l}}=\sum_{i=1}^{r}\frac{1}{n_{i}}\sum_{l=1}^{n_{i}-1}\frac{\sum_{\alpha=0}^{n_{i}-1}N_{i,\alpha}\cdot\zeta_{n_{i}}^{l(\alpha+\nu_{h_{i}^{l}})}}{1-\zeta_{n_{i}}^{l}}.$
@@
We want to point out that in order to determine the integers $N_{i,\alpha}$ it is not necessary to know the representation $\varrho\colon G\to\operatorname{GL}(V)$. Indeed, it suffices to know the character $\chi\in\operatorname{Irr}(G)$ of $\varrho$, as explained in the “Computational Remark 2.9” of [FG16], where a proof of the Chevalley-Weil formula for $m=1$ is included too.
@@ 3. Pluricanonical Systems of an Abelian Cover
Let $X\cong(C_{1}\times\cdots\times C_{n})/G$ be a variety isogenous to a product of unmixed type. Consider the associated cover (see diagram (2.1))
@@
In the preceding section, we described how to determine the character $\chi_{m}$ of the representation
@@
Let $\pi\colon X\to Y$ be a finite abelian cover between complex manifolds with Galois group $G$, and let $m$ be a non negative integer. The $G$-action on $\pi_{*}(K_{X}^{\otimes m})$ via pullback of holomorphic differential pluriforms yields the eigensheaf-decomposition
$\pi_{*}(K_{X}^{\otimes m})=\bigoplus_{\chi\in\operatorname{Irr}(G)}\pi_{*}(K_{X}^{\otimes m})^{\chi}\,$,
where, for an open set $U\subseteq Y$,
$\pi_{\ast}(K_{X}^{\otimes m})^{\chi}(U):=\{\omega\in K_{X}^{\otimes m}(\pi^{-1}(U))\ \mid\ g^{\ast}\omega=\chi(g)\cdot\omega\quad\makebox{for all}\quad g\in G\}.$
Given a character $\chi\in\operatorname{Irr}(G)$}, our goal is to provide an explicit description of the line bundle $\pi_{*}(K_{X}^{\otimes m})^{\chi}$.
@@
$a_{j}(x)t^{k-m}\frac{(dx_{1}\wedge\ldots\wedge dx_{n})^{\otimes m}}{t^{(d-1)m}}=\psi_{ij}^{\ast}\bigg(a_{i}(x)t^{k-m}\frac{(dx_{1}\wedge\ldots\wedge dx_{n})^{\otimes m}}{t^{(d-1)m}}\bigg)=a_{i}(x)g_{ij}(x)^{k-dm}t^{k-m}\frac{(dx_{1}\wedge\ldots\wedge dx_{n})^{\otimes m}}{t^{(d-1)m}}.$
This implies $a_{i}(x)=g_{ij}(x)^{dm-k}a_{j}(x)$. In other words, the collection $\{a_{i}(x)\}_{i}$ defines a section of the line bundle $L^{\otimes(dm-k)}=\mathcal{O}_{Y}(mB)\otimes L^{\otimes(-k)}$. It follows that the map
@@
(3.1) $L_{\chi_{k}}\cong L^{\otimes k},\qquad 0\leq k\leq d-1.$
@@
$\pi_{\ast}(K_{X}^{{\otimes m}})^{\chi_{k}}=\pi_{\ast}(K_{X}^{{\otimes m}})^{\chi_{(k+ld)}}\cong L^{\otimes(dm-k-ld)}\otimes K_{Y}^{{\otimes m}}=L^{\otimes(dm-k-ld-d+k)}\otimes L^{\otimes(d-k)}\otimes K_{Y}^{{\otimes m}},$
where the isomorphism $L^{\otimes(dm-k-ld)}\otimes K_{Y}^{{\otimes m}}\to\pi_{\ast}(K_{X}^{{\otimes m}})^{\chi_{(k+ld)}}$ is given by
$s\mapsto t^{r}\otimes\pi^{*}s\,,\qquad r:=k+ld-m.$
If $k\neq 0$, by (3.1) of Remark 3.4 the previous chain of equalities amounts to
@@
Then the formula follows from Remark 3.4. ∎
Let us now deal with the general case. Given a finite abelian cover $\pi\colon X\to Y$ between complex manifolds with Galois group $G$, we denote by $R$ the ramification divisor and by $D$ the branch divisor. If $T$ is an irreducible component of $R$ and $p\in T$ a general point, we denote by $H$ the stabilizer group of $p$, which in fact does not depend on the choice of $p$, and hence it is also called the stabilizer of $T$. Then, by Cartan’s lemma [Cartan, Lemma 1], the tangent representation at $p$ is faithful and looks as follows
@@
(3.2) $p^{\ast}(\mathcal{O}_{Y}(D_{(H,\psi)}))=\mathcal{O}_{Z}(B)\qquad\makebox{and}\qquad p^{\ast}K_{Y}^{\otimes m}=K_{Z}^{\otimes m}.$
@@
(3.3) $p_{\ast}M^{\otimes(-k)}=\bigoplus_{\underset{\eta_{|H}=\psi^{k}}{\eta\in\operatorname{Irr}(G)}}L^{-1}_{\eta}\,,\qquad\text{ where }\quad L^{-1}_{\eta}=(\pi_{*}\mathcal{O}_{X})^{\eta}.$
@@
(3.4) $(\pi_{\ast}K_{X}^{\otimes m})^{\chi}=(p_{\ast}g_{\ast}K_{X}^{\otimes m})^{\chi}\cong\left(\bigoplus_{k=0}^{d-1}p_{\ast}\left(\mathcal{O}_{Z}\left(\mu_{k}B\right)\otimes M_{\psi^{k}}^{-1}\otimes K_{Z}^{\otimes m}\right)\right)^{\chi}$
@@
By (3.2) each summand in (3.4) is isomorphic to
$\left(p_{\ast}\left(p^{\ast}\left(\mathcal{O}_{Y}\left(\mu_{k}D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\right)\otimes M_{\psi^{k}}^{-1}\right)\right)^{\chi}\cong\mathcal{O}_{Y}\left(\mu_{k}D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\otimes\left(p_{\ast}M_{\psi^{k}}^{-1}\right)^{\chi},$
where the isomorphism is given by projection formula. By using (3.3) we obtain
$(\pi_{\ast}K_{X}^{\otimes m})^{\chi}\cong\bigoplus_{k=0}^{d-1}\left(\mathcal{O}_{Y}\left(\mu_{k}D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\otimes\Bigg(\bigoplus_{\underset{\eta_{|H}=\psi^{k}}{\eta\in\operatorname{Irr}(G)}}L^{-1}_{\eta}\Bigg)^{\chi}\ \right)\,.$
Since there exists a unique integer $0\leq k_{0}\leq d-1$ such that $\chi_{|H}=\psi^{k_{0}}$, only one summand survives:
$(\pi_{\ast}K_{X}^{\otimes m})^{\chi}\cong\mathcal{O}_{Y}\left(\mu_{k_{0}}D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\otimes L^{-1}_{\chi}.$
Following the maps, we get the isomorphism of line bundles
@@
$r:=k_{0}-m+\left\lceil\frac{m-k_{0}}{d}\right\rceil d,$
by the isomorphism given in the proof of Lemma 3.5.
@@
(3.5) $\pi_{\ast}(K_{X}^{\otimes m})^{\chi}\cong\mathcal{O}_{Y}\left(\sum_{(H,\psi)}\mu(m,H,\psi,\chi)\ D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\otimes L_{\chi}^{-1},$
@@
(3.6) $H^{0}(X,K_{X}^{\otimes m})=\bigoplus_{\chi\in\operatorname{Irr}(G)}\prod_{(H,\psi)}t_{(H,\psi)}^{r(m,H,\psi,\chi)}\ \pi^{\ast}H^{0}\left(Y,\mathcal{O}_{Y}\left(\sum_{(H,\psi)}\mu(m,H,\psi,\chi)\ D_{(H,\psi)}\right)\otimes K_{Y}^{\otimes m}\otimes L_{\chi}^{-1}\right),$
@@
Since $\pi_{\ast}(K_{X}^{\otimes m})$ is reflexiv by [Har80, Corollary 1.7], it is enough to prove the result when $X$ is smooth. Our discussion above shows that the line bundle isomorphism in (3.5) holds in codimension 1. Hence, it holds globally by Hartog’s extension theorem and the statement follows immediately. ∎
@@
Formula (3.5) in Theorem 3.7 has been derived by Pardini [Par91, Proposition 4.1, c)] for $m=1$ and by Bauer-Pignatelli [BP21, Proposition 2.5] for $m=2$. They write them as follows
@@
In [Lie03] Liedtke derived formula (3.6) in Theorem 3.7 for $m=1$ and wrote it as follows:
@@
(3.7) $L_{\chi^{-1}}=L_{\chi}^{-1}\otimes\mathcal{O}_{Y}\bigg(\sum_{\underset{\chi_{|H}\neq\psi^{0}}{(H,\psi)}}D_{(H,\psi)}\bigg).$
Hence, by using (3.7) together with (3.5), we obtain
@@
whence we obtain Bauer-Pignatelli’s result. Assume now $m=1$. By using (3.7) together with (3.6), where we replace $\chi^{-1}$ with $\chi$, we get
@@ 3.1. Pluricanonical Systems of an Unmixed VIP with Abelian Group
Let $X\cong(C_{1}\times\dots\times C_{n})/G$ be a variety isogenous to a product of unmixed type with abelian group $G$. Recall that we have the diagram (2.1):
${C_{1}\times\dots\times C_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f_{1},\dots,f_{n})}$$\scriptstyle{\Xi}$${C_{1}^{\prime}\times\dots\times C_{n}^{\prime}}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$
@@
In summary, this provides a method to compute each summand in the decomposition (3.8).
@@
Consequently, the integers $r(m,H,\psi,\chi)$ and $\mu(m,H,\psi,\chi)$ are determined immediately. The only remaining data needed to apply Theorem 3.7 are the line bundles $L_{\chi}$, for which the following formula holds (see [Par91, (2.14) of Proposition 2.1]):
@@
where $d_{\chi}:=\operatorname{ord}(\chi)$. Thus, we have a concrete procedure to compute $L_{\chi}^{\otimes d_{\chi}}$, which determines the line bundle $L_{\chi}$ up to torsion. In the special case where all curves $C_{i}^{\prime}=\mathbb{P}^{1}$ we can immediately recover $L_{\chi}$ itself.
4. Pluricanonical Maps of an Abelian Cover
Our goal is to study the birationality of the $m$-canonical map ($m\geq 1$) of an unmixed VIP with abelian Galois group. For this purpose, we consider an arbitrary abelian cover $\pi\colon X\to Y$ and recall two birationality criteria established by Catanese in [Cat18] for the canonical map of $X$. We present them here in the generalized setting of the $m$-canonical map